The National Climate Data Center reports that the mean annual heating degree-days for the period 1949–2006 was 4500 with a standard deviation of 200. Assume that the shape of the data distribution is not known.
At least ___ percent of years have heating degree-days between 4100 and 4900.

Answer by ewatrrr(24785)

z[4900] = (4900 - 4500)/200 = 2
z(4100) = (4100-4500) /200 = -2
two standard deviations from the mean account for about 95%